Which type of unemployment is a permanent and normal phenomenon in the economy? frictional unavoidable structural...

  1. Which type of unemployment is a permanent and normal phenomenon in the economy?

    frictional

    unavoidable

    structural

    cyclical

2 points   

QUESTION 22

  1. Market Basket 1990 (Base Year) 2010 2011
    Product Quantity Price Price Price
    A 8 $4 $6 $7
    B 12 $6 $9 $9
    C 21 $7 $9 $10
    D 16 $10 $14 $13

    If your answer is not a whole number, round to two decimal places.

    What is the cost of the market basket in 1990? $

    What is the cost of the market basket in 2010? $

    What is the cost of the market basket in 2011? $

    What is CPI in 1990?

    What is CPI in 2010?

    What is CPI in 2011?

    What was the inflation rate in 2011? %
Answer #1

Answer)

1)Which type of unemployment is a permanent and normal phenomenon in the economy?

cyclical

Answer 22)

Cost of basket in 1990 = 8*4+12*6+21*7+16*10

=411

Cost of basket in 2010 = 8*6+12*9+21*9+16*14

=569

Cost of basket in 2011 = 8*7+12*9+21*10+16*13

=582

CPI = cost of basket in current period /cost of basket in base period * 100

CPI in 1990 = 411/411 * 100 = 100

CPI in 2010 = 569/411 * 100 = 138.44

CPI in 2011 = 582/411 * 100 = 141.60

Inflation rate in 2011 = CPI (2011) - CPI(2010) /CPI(2010) * 100

= 141.60 - 138.44/138.44 * 100

= 2.28%

If you have any doubts please comment...